OpenStudy (anonymous):

Calculate the volume occupied (a) 16 g of O2. (b) 1.5 mole of N2. (c) 6.02 * 10^20 molecules.

OpenStudy (anonymous):

Volume occupied?? Level of difficultly increasing for me with each your question.. Ha ha ha..

OpenStudy (anonymous):

:P .. so we'll both study together !

OpenStudy (anonymous):

1 mole of 02 = 22.4 L

OpenStudy (anonymous):

Would we use this concept here ??

OpenStudy (anonymous):

okay .. is this like one mole = 6.023810^23 molecules . mole of an atom or compound = 22.4 L ?

OpenStudy (anonymous):

lol .. you asking me ?

OpenStudy (anonymous):

No I am not asking.. Ha ha ha..

OpenStudy (anonymous):

then ?

OpenStudy (anonymous):

Use this: 1 mole of 02 = 22.4 L 32g = 22.4 find for 16 g..

OpenStudy (anonymous):

Again 16g is half of 32g remember..

OpenStudy (anonymous):

okay .. so 22.4 * 16 / 32 = 358.4/32 = 11.2 L correct ?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

Do the same for b part..

OpenStudy (anonymous):

okay !

OpenStudy (anonymous):

Do not convert that into grams..

OpenStudy (anonymous):